OpenStudy (anonymous):

Write the following as an algebriac expression in "x" cot(arccos(3x)), 0 ≤ x ≤ 1/3

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

This problem makes my head hurt :( lol

OpenStudy (misty1212):

this is really not that bad at all first step: draw a triangle with some angle \(\theta\) labelled next step, since \[\cos(\theta)=3x\] put a \(3x\) on the "adjacent" side and make the hypotenuse 1

OpenStudy (misty1212):



OpenStudy (misty1212):

there is a picture of an angle whose cosine is \(3x\) what you need is the "opposite" side which you find via pythagoras

OpenStudy (misty1212):

let me know what you get it should have a square root in it

OpenStudy (anonymous):

:o let me try

OpenStudy (anonymous):

sqrt(-9x^2 + 1) ?

OpenStudy (misty1212):

yes although \[\sqrt{1-9x^2}\] looks neater

OpenStudy (anonymous):

haha alright

OpenStudy (misty1212):

now you want the cotangent of that angle, which is "adjacent over opposite" so you are pretty much done, just write it i

OpenStudy (anonymous):

Yeah I got it form here, you just made that so easy it's ridiculous. Thank you so much

OpenStudy (misty1212):

it is ridiculously easy if you know what you are doing hope your head does not hurt anymore \[\color\magenta\heartsuit\]

OpenStudy (anonymous):

It's alleviated :)